Flora11:

The average weight of two boys is 48 kg. If one boy is 6 kg heavier than the other, find the height of the heavier boy.

ramen:

we know that there are 2 boys and their average is 48. based off this info, since there are 2 boys, we can work backwards. 48*2=96

ramen:

what two numbers add up to 96 but have a difference of 6?

ramen:

i did 48*2 because with averages, since there are two boys, we would originally divide by 2 to find the average of the two numbers. we need the number that would divide by 2 and end up being 48. thats where 96 comes in.

justjm:

You could do it the way @ramen told you or you could use a more prescriptive method: Let us denote the weight of the first boy as x, and second by as y. \[W_{ave}=\frac{x+y}{2}\] \[∴ 1)...\frac{x+y}{2}=48~kg\] You are also given that one boy is 6 kg heavier than the other (i.e., x is 6 kg heavier than y). \[2)...x=6+y\] Now you can employ substitution to find the values of x and y, and your answer would be the weight (I believe the question saying 'height' is an error) of the heavier boy (i.e., weight of x)

TheSmartOne:

how tall is the bigger boi tho